# Imagine Beyond! Distributionally Robust Auto-Encoding for State Space Coverage in Online Reinforcement Learning

### v2

@@ 2.1 Problem Statement: Unsupervised Goal-conditioned Reinforcement Learning
In this work, we consider the multi-goal reinforcement learning (RL) setting, defined as an extended Markov Decision Process (MDP) ${\cal M}=<S,T,A,G,R_{g},S_{0},X>$, where $S$ is a set of continuous states, $T$ is the transition function, $A$ the set of actions, $S_{0}$ the distribution of starting states, $X$ the observation function and $R_{g}$ the reward function parametrized by a goal $g\in G$. In our unsupervised setting, we are interested in finding control policies that are able to reach any state in $S$ from the distribution of starting states $S_{0}$. Thus, we consider that $G\equiv S$. $S$ being continuous, we set the reward function $R_{g}$ as depending on a threshold distance $\delta$ from the goal $g$: for any state $s\in S$, we consider the sparse reward function $R_{g}(s)={\mathbbm{1}}[||s-g||_{2}<\delta]$. $R_{g}(s)>0$ is only possible once per trajectory. For simplicity, we also consider that any pixel observation $x\in X$ corresponds to a single state $s\in S$. Thus, given an horizon $T$, the optimal policy is defined as $\pi^{*}=\max_{\pi}\mathbb{E}_{g\in G}\mathbb{E}_{\tau\sim\pi(\tau|g)}[\sum_{t=0}^{T}R_{g}(s_{t})]$, where $\tau=(s_{0},a_{1},s_{1},...,s_{T})$ is a trajectory, and $\pi(\tau|g)$ is the distribution of trajectories given $g$ in the MDP when following policy $\pi$.
@@ 2.4 Distributionally Robust Optimization
Assuming $\mathcal{Q}$ as a set of distributions that are absolutely continuous with respect to $p$ [^1] In the situation where all distributions in ${\cal Q}$ are absolutely continuous with respect to $p$, for all measurable subset $A\subset X\times Y$ and all $q\in\mathcal{Q}$, $q(A)>0$ only if $p(A)>0$., the inner maximization problem of DRO can be reformulated using importance weights $r(x,y)$ such that $q=rp$ (Michel et al., 2022). In that case, we have:
@@
[…]
@@ 3.1 DRO-VAE
[…]
@@
where the only difference is that the inner maximization considers an approximation ~$\tilde{{\cal L}}_{\theta,\phi}(x)\approx\log{p_{\theta,\phi}(x)}$. ~$\tilde{{\cal L}}_{\theta,\phi}(x)$ is estimated via Monte-Carlo importance sampling, as ~$\tilde{{\cal L}}_{\theta,\phi}(x)=\log\sum_{j=1}^{M}\exp(\log p_{\theta}(x|z^{j})+\log p_{\theta}(z^{j})-\log q_{\phi}(z^{j}|x))-\log(M)$ given $M$ samples $z^{j}$ from $q_{\phi}(z^{j}|x)$ for any $x$, which can be computed accurately (without loss of low log values) using the LogSumExp trick.
@@ 4 Experiments
Our experiments seek to highlight the impact of drag on the efficiency of GCRL from pixel input. As depicted in Figure 2, we structure this section around two experimental steps that seek to answer the two following research questions in isolation:
@@
We consider two kinds of environments, with observations as images of size $82\times 82$. In Pixel continuous PointMazes, we evaluate the different algorithms over 4 hard-to-explore continuous point mazes. The action space is a continuous vector $(\delta x,\delta y)=[0,1]^{2}$. Episodes start at the bottom left corner of the maze. Reaching the farthest area requires at least 40 steps in any maze. States and goals are pixel top-down view of the maze with a red dot highlighting the corresponding xy position. Pixel Reach-Hard-Walls is adapted from the Reach-v2 MetaWorld benchmark (Yu et al., 2021). We add 4 brick walls that limit the robotic arm’s ability to move freely. At the start of every episode, the robotic arm is stuck between the walls.
@@ 4.1 Representation Learning strategy
Results in Figure 3 show the evolution of the success coverage over 4M steps. We see that drag significantly outperforms rig and Skew-Fit. These results corroborate that online representation learning with rig is unable to overcome an exploration bottleneck. Therefore, a rig agent can only explore and control a very small part of the environment. Furthermore, the success coverage of rig is systematically capped to a certain value. Skew-Fit is often able to overcome the exploration bottleneck but suffers from high instability, which indeed corresponds to the main drawback of non-parametric DRO, highlighted in Michel et al. (2021). Therefore, Skew-Fit is unable to reliably maximize the success coverage. On the other hand, drag is more stable due to the use of parametric likelihood ratios and is able to maximize the success coverage. Additional results and visualizations on these experiments are presented in Appendix D.1. In particular, they show a greatly better organized latent space with drag than with other approaches.
@@ A.3 Methods Hyper-parameters
Optimization interval (in agent steps) $freqOpt$ 10e3
Nb of training steps $nbEpochs$ 50
Size of training buffer $|{\cal R}|$ 1e6
Number of samples per epoch $N$ 1e3
@@
A.4 Compute ressources & code assets acknowledgments
This work was performed with 35,000 GPU hours on NVIDIA V100 GPUs (including main experiments and ablations).
@@ Appendix B DRAG algorithm
1:  Input: a GCP $\pi_{\theta}$, a VAE: encoder $q_{\phi}(z|x)$ and smoothly updated version $q_{\hat{\phi}}(z|x)$, decoder $p_{\phi}(x|z)$, latent prior $p(z)$, DRO Neural Weighter $r_{\psi}$, buffers of transitions ${\cal B}$, reached states ${\cal R}$, train size $N$, batch-size $n$, number $m$ of sampled noises for each VAE training input, number $M$ of Monte Carlo samples used to estimate $\tilde{L}_{\theta,\phi}$ for each input image, temperature $\lambda$, number of optimization epochs $nEpochs$, frequence of VAE and policy optimization $freqOpt$.
@@
7:     for $nEpochs$ epochs do
8:        Sample a train set of N states $\Gamma$ from ${\cal R}$
@@
15:                  Update encoder $q_{\phi}$ and decoder $p_{\phi}$ by one step of Adam optimizer on $-{\cal L}^{\text{\tiny VAE-DRO}}_{\theta,\phi,r_{\psi}}(\{x_{i}\}_{i=1}^{n})$, as defined in (5), with $m$ sampling noises $(\epsilon_{i}^{j})_{j=1}^{m}$ for each $x_{i}$.
@@
21:            Improve agent with any Off-Policy RL algorithm (e.g., tqc, ddpg, sac…) using transitions from $\cal B$;
@@ C.1 Non-parametric solution of DRO
Following the Karush-Kuhn-Tucker conditions applied to the derivative of the Lagrangian function ${\cal L}$ of this problem in $r_{i}$ for any given $i\in[[1,N]]$, we obtain:
@@ C.2 Application to GCRL with VAE and Relation to Skew-Fit
where $p(z)$ is the prior over latent encodings of the data $x$, $p_{\theta}(x|z)$ is the likelihood of $x$ knowing $z$ and $q_{\phi}(z|x)$ the encoding distribution of data points. As stated in Section 3.1, this can be estimated on a set of $m$ samples for each data point using the log-approximator: ~$\tilde{{\cal L}}_{\theta,\phi}(x)=\log\sum_{j=1}^{M}\exp(\log p_{\theta}(x|z^{j})+\log p_{\theta}(z^{j})-\log q_{\phi}(z^{j}|x))-\log(M)$.
@@
for any $\alpha<0$. Setting $\alpha=-\frac{1}{\lambda}$, we get $p_{\text{skewed}}(\mathbf{x_{i}})\propto e^{-\tilde{{\cal L}}_{\theta,\phi}(x_{i})/\lambda}$, for any temperature $\lambda>0$. Reusing the result from Section C.1, this is fully equivalent to the analytical closed-form solution of DRO when applied to ~$-\tilde{{\cal L}}_{\theta,\phi}(x_{i})$ as we use in our DRO-VAE approach. Using $p_{skewed}$ with $\alpha=-\frac{1}{\lambda}$ for weighting training points of a VAE thus exactly corresponds to the non-parametric version our drag algorithm.
@@ D.1 Visualization of Learned Latent Representations
In order to gain a deeper insight into the performance of rig, Skew-Fit, and drag, we show in Figure 9 and Figure 9 the parallel evolution of the prior sampling $z\sim\mathcal{N}(0,I)$, and the corresponding learned 2D representations for the maze environment in Figure 9.
@@
D.2 Ablations
@@ D.2.1 Impact of $\lambda$
D.2.2 Impact of $M$ (number of samples for the estimation of $\tilde{L}_{\theta,\phi}$ )
@@
D.3 RIG+Goal selection criterion
@@ Appendix E Limitations
While our study makes progress on learning representations online and generating intrinsic goals from high-dimensional observations, it does not address how to measure when a goal has truly been achieved. Throughout our experiments, we relied on a simple sparse reward $r_{t}={\mathbbm{1}}\big{[}||z_{x_{t}}-z_{g}||_{2}<\delta\big{]}$ which, although common in goal-conditioned RL, sidesteps the challenges of defining a dense feedback signal. In particular, the Euclidean metric used in dense rewards often fails to reflect the true topology of the environment and can mislead the agent.

### v3

@@ 2.1 Problem Statement: Unsupervised Goal-conditioned Reinforcement Learning
[…]
@@ 2.4 Distributionally Robust Optimization
Assuming $\mathcal{Q}$ as a set of distributions that are absolutely continuous with respect to pp [^1] In the situation where all distributions in $\cal Q$} are absolutely continuous with respect to $p$, for all measurable subset $A\subset X\times Y$ and all $q\in\mathcal{Q}$, $q(A)>0$ only if $p(A)>0$., the inner maximization problem of DRO can be reformulated using importance weights $r(x,y)$ such that $q=rp$ (Michel et al., 2022). In that case, we have:
@@
where the constraint ensures that the $q$ function keeps a valid integration property for a distribution (i.e., $\int_{{\cal X},{\cal Y}}q(x,y)dxdy=1$). The term $\lambda\log r$ is a relaxation of a KL constraint, which ensures that $q$ does not diverge too far from pp [^2] This can be seen easily, observing that: $KL(q||p)=\int q(x)\log(q(x)/p(x))dx=\int p(x)r(x)\log r(x)dx$).. $\lambda$ is an hyper-parameter that acts as a regularizer ensuring a trade-off between generalization to shifts (low $\lambda$) and accuracy on training distribution (high $\lambda$).
@@ 3.1 DRO-VAE
Classic VAE learning aims at minimizing the negative log-likelihood: ${\cal L}=-\mathbb{E}_{x\sim p(x)}\log p_{\theta,\phi}(x)$, with $p_{\theta,\phi(x)}$ the predictive posterior, which can be written as: $p_{\theta,\phi}(x)=\int p(z)p_{\theta}(x|z)dz$, where $p(z)$ is a prior over latent encoding of the data $x$, commonly taken as ${\cal N}(0,I)$, and $p_{\theta}(x|z)$ is the likelihood of $x$ knowing $z$ and the parameters of the decoding model $\theta$. Given that this marginalization can be subject to very high variance, the idea is to use an encoding distribution $q_{\phi}(z|x)$ to estimate this generation probability (Kingma and Welling, 2013). For any distribution $q_{\phi}$ such that $q_{\phi}(z|x)>0$ for any $z$ with $p(z)>0$, we have: $p_{\theta,\phi}(x)=\mathbb{E}_{z\sim q_{\phi}(z|x)}p(z)p_{\theta}(x|z)/q_{\phi}(z|x)$.
@@
where the only difference is that the inner maximization considers an approximation $\tilde{{\cal L}}_{\theta,\phi}(x)\approx\log{p_{\theta,\phi}(x)}$. $\tilde{{\cal L}}_{\theta,\phi}(x)$ is estimated via Monte-Carlo importance sampling, as $\tilde{{\cal L}}_{\theta,\phi}(x)=\log\sum_{j=1}^{M}\exp(\log p_{\theta}(x|z^{j})+\log p_{\theta}(z^{j})-\log q_{\phi}(z^{j}|x))-\log(M)$ given $M$ samples $z^{j}$ from $q_{\phi}(z^{j}|x)$ for any $x$, which can be computed accurately (without loss of low log values) using the LogSumExp trick.
@@ 4 Experiments
Our experiments seek to highlight the impact of drag on the efficiency of GCRL from pixel input [^4] The code is available at https://github.com/nicolascastanet/DRAG. As depicted in Figure 2, we structure this section around two experimental steps that seek to answer the two following research questions in isolation:
@@
We consider two kinds of environments, with observations as images of size $82\times 82$. Additional results on image of size $128\times 128$ are also provided in appendix D.5. In Pixel continuous PointMazes, we evaluate the different algorithms over 4 hard-to-explore continuous point mazes. The action space is a continuous vector $(\delta x,\delta y)=[0,1]^{2}$. Episodes start at the bottom left corner of the maze. Reaching the farthest area requires at least 40 steps in any maze. States and goals are pixel top-down view of the maze with a red dot highlighting the corresponding xy position. Pixel Reach-Hard-Walls is adapted from the Reach-v2 MetaWorld benchmark (Yu et al., 2021). We add 4 brick walls that limit the robotic arm’s ability to move freely. At the start of every episode, the robotic arm is stuck between the walls.
@@ 4.1 Representation Learning strategy
Results in Figure 3 show the evolution of the success coverage over 4M steps. We see that drag significantly outperforms rig and Skew-Fit. These results corroborate that online representation learning with rig is unable to overcome an exploration bottleneck. Therefore, a rig agent can only explore and control a very small part of the environment. Furthermore, the success coverage of rig is systematically capped to a certain value. Skew-Fit is often able to overcome the exploration bottleneck but suffers from high instability, which indeed corresponds to the main drawback of non-parametric DRO, highlighted in Michel et al. (2021). Therefore, Skew-Fit is unable to reliably maximize the success coverage. On the other hand, drag is more stable due to the use of parametric likelihood ratios and is able to maximize the success coverage. Additional results and visualizations on these experiments are presented in Appendix D.1. In particular, they show a greatly better organized latent space with drag than with other approaches. We also show in appendix D.2 that drag obtains better latent representations in terms of the trustworthiness score (Venna and Kaski, 2001), which measures the preservation of the local neighborhood structure in the input space. Metrics regarding computational runtime are reported in appendix D.6.
@@ A.3 Methods Hyper-parameters
Optimization interval (in agent steps) freqOptfreqOpt 10e3
Nb of training steps nbEpochsnbEpochs 50
[…]
Number of samples per epoch NN 1e3
@@
A.4 Compute ressources & code assets
This work was performed with 35,000 GPU hours on NVIDIA V100 GPUs (including main experi- ments and ablations).
@@ Appendix B DRAG algorithm
[…]
@@
7:  for nEpochsnEpochs epochs do
[…]
@@
15:       Update encoder $q_{\phi}$ and decoder $p_{\phi}$ by one step of Adam optimizer on ${\cal L}^{\text{\tiny VAE-DRO}}_{\theta,\phi,r_{\psi}}(\{x_{i}\}_{i=1}^{n})$, as defined in (5), with $m$ sampling noises $(\epsilon_{i}^{j})_{j=1}^{m}$ for each $x_{i}$.
@@
[…]
@@ C.1 Non-parametric solution of DRO
[…]
@@ C.2 Application to GCRL with VAE and Relation to Skew-Fit
where $p(z)$ is the prior over latent encodings of the data $x$, $p_{\theta}(x|z)$ is the likelihood of $x$ knowing $z$ and $q_{\phi}(z|x)$ the encoding distribution of data points. As stated in Section 3.1, this can be estimated on a set of $m$ samples for each data point using the log-approximator: $\tilde{{\cal L}}_{\theta,\phi}(x)=\log\sum_{j=1}^{M}\exp(\log p_{\theta}(x|z^{j})+\log p_{\theta}(z^{j})-\log q_{\phi}(z^{j}|x))-\log(M)$.
@@
for any $\alpha<0$. Setting $\alpha=-\frac{1}{\lambda}$, we get $p_{\text{skewed}}(\mathbf{x_{i}})\propto e^{-\tilde{{\cal L}}_{\theta,\phi}(x_{i})/\lambda}$, for any temperature $\lambda>0$. Reusing the result from Section C.1, this is fully equivalent to the analytical closed-form solution of DRO when applied to $\tilde{{\cal L}}_{\theta,\phi}(x_{i})$ as we use in our DRO-VAE approach. Using $p_{skewed}$ with $\alpha=-\frac{1}{\lambda}$ for weighting training points of a VAE thus exactly corresponds to the non-parametric version our drag algorithm.
@@ D.1 Visualization of Learned Latent Representations
In order to gain a deeper insight into the performance of rig, Skew-Fit, and drag, we show in Figure 9 the parallel evolution of the prior sampling $z\sim\mathcal{N}(0,I)$, and the corresponding learned 2D representations for the maze environment.
@@
D.2 Representation quality
We assess the quality of the obtained representations in term of the trustworthiness score [Venna and Kaski, 2001], which measures to what extent the local neighborhood structure in the input space is preserved in the latent space. Higher trustworthiness indicates better preservation of task-relevant information across the encoding. We computed this trustworthiness metric (with 5 nearest neighbors) using a batch of 1K uniformly sampled observations from the valid state space, identical to the batch used for success coverage. Our experiments show that DRAG consistently improves this score over baseline encoders across environments.
Table 4: Mean trustworthiness of obtained representations across environments after 4M training steps.
Methods Trustworthiness [Venna & Kaski, 2001]
DRAG 98.7%
SKEWFIT 93.3%
RIG 88.5%
D.3 Ablations
@@ D.3.1 Impact of $\lambda$
D.3.2 Impact of $M$ (number of samples for the estimation of $\tilde{L}_{\theta,\phi}$ )
@@
D.4 RIG+Goal selection criterion
@@
D.5 Image size study
To complement our experiments, we ran the three methods (i.e. rig, Skew-Fit and drag) on all environments with a higher image resolution (128x128 instead of 82x82 in the main experiments). We observe that all methods degrade slightly with higher resolution, but drag retains a clear advantage.
Table 5: Mean success coverage across maze and robotic environment for 128x128 pixels observations (success coverage for 82x82 is given for comparison).
Methods 1M 2M 3M 4M steps
DRAG 42% 58% 71% 79%
    82x82 comparison 46% 66% 74% 81%
SKEWFIT 28% 44% 57% 65%
    82x82 comparison 33% 51% 59% 66%
RIG 22% 29% 35% 38%
    82x82 comparison 28% 34% 39% 42%
D.6 Runtime
To assess the impact of the overhead induced by the additional likelihood estimation (for SKEWFIT and DRAG compared to RIG) and the use of a neural weighter (for DRAG), Table 6 reports mean execution runtime for Maze and Metaworld environments averaged over 6 seeds on each environment. We observe that DRAG and SKEWFIT require on average about 200 seconds for 20k steps, against 176 seconds for RIG. This slight overhead is negligible compared to the time spent in environment interactions. This is because we only use one sample (i.e. ) for the likelihood estimation in DRAG and SKEWFIT (see appendix D.3.2 for a comparative study with more samples, which does not impact the results of both methods). Also, no notable difference in runtime is observed when comparing DRAG to SKEWFIT.
Table 6: Mean execution runtime on V100 GPU (32GB) for MAZE and METAWORLD environments.
Method Time per 20K steps (s) Time for 4M steps (s)
DRAG 211 42,200 (11.72 h)
SKEWFIT 208 41,600 (11.55 h)
RIG 176 35,200 (9.77 h)
To better link performance to runtime, Table 7 also shows success coverage over wall-clock time (82×82, same runs as Fig. 3):
Table 7: Success coverage metric per hour.
Hour DRAG (%) SKEWFIT (%) RIG (%)
1 34 31 26
3 45 42 28
5 60 50 32
7 70 58 36
9 77 63 40
11 81 66 42
@@ Appendix E Limitations
[…]
@@
Notably, contrastive methods [Stooke et al., 2021] and Forward-Backward approaches [Touati and Ollivier, 2021] aim to incorporate dynamics by bringing temporally adjacent states closer or by modeling universal rewards. However, these methods generally assume access to transitions from a representative part of the environment. To our knowledge, they do not include any explicit mechanism to avoid collapse onto narrow parts of the state space—a critical issue in hard exploration settings without expert priors.
Addressing this limitation is precisely the aim of our DRO-based reweighting, which promotes state space coverage even in sparse reward regimes. While our implementation focuses on $\beta$-VAEs for interpretability and disentanglement, our DRO framework is general and not tied to a specific representation architecture.
In fact, any representation learning method trained from replay buffer samples using a loss $\ell(f_{\theta}(x))$ can be reweighted using the DRO objective:
$\min_{\theta}\max_{r}\frac{1}{N}\sum_{i=1}^{N}r(x_{i})(\ell(f_{\theta}(x_{i})))-\lambda r(x_{i})\log r(x_{i})\qquad\text{s.t. }\frac{1}{N}\sum_{i=1}^{N}r(x_{i})=1$
This includes diffusion-based decoders, contrastive losses (e.g., InfoNCE), and temporal-difference-based objectives (e.g., in Forward-Backward RL). For generative losses, as in VAEs or diffusion models, the optimization must rely on likelihood estimates (i.e., $\log p_{\theta,\phi}(x)$ ), similarly to our alternate optimization strategy described in (4).